Find the height over the earth’s surface at which the weight of a body becomes half of its value at the surface.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Let the height be h ∴(1/2) GM/R^2 = GM/(R + h)^2 Or 2R^2 = (R + h)^2 Or √2 R = R = R + h Or h = (r base 2 - 1)R